Friday, February 26, 2010

Web Hosting Account Server service restored.

The web hosting server has been brought back up.

Web Hosting Account Server Issue

The web hosting account server (people/faculty/ experiencing a network issue. The Web Services and Network Team are working to resolve the issue.

E-mail flowing again

Power issue for several boxes that held up e-mail.  All queued messages are being delivered.


Email Issue as of 1:41 am

Incoming email appears to have stop flowing in from 1:41 am this morning. ITS is investigating.

Thursday, February 25, 2010

Blackboard Maintenance - Service Outage Monday 03/01 7:30 a.m.

For technical reasons, we have scheduled Blackboard hardware maintenance which will cause service outage. Blackboard service will be shut down Monday, March 1 at 7:30 a.m. Estimated time needed to accomplish necessary service/repair is 30 minutes. Please log out and finish your work BEFORE 7:30 a.m.; you may experience loss of data if you don't do so. Update on status of Blackboard service will be announced in this blog.
Sorry for inconvenience that this operation may cause to you and thank you for understanding,


Weblink is Available Now

Apecs and Weblink are available. Weblink involved more steps than expected but is functioning fine now. Thank you, ITS dept.

UNISYS mainframe required reboot: Apecs and Weblink will be up in about 5 minutes

The UNISYS mainframe experienced a lockup and required a reboot. The APECS and Weblink environment will be available in approximately 5 minutes. ITS Dept.

Monday, February 22, 2010

All systems operational

All systems are operational at this time.


Thursday, February 18, 2010

TC Portal Restart 2/19 @ 4:00 am


The TC portal will be stopped and started for some maintenance work around 4:00 am and lasting 30 minutes.  If you are on at this time, please log off the system.  Active connections at 4:00 am may be abruptly ended.

Sunday, February 14, 2010

Resolved - Internet Provider Issue

The issue we were experiencing with out internet provider has been resolved.

-ITS Network Group

Internet Provider Issue

We just discovered that we are experiencing a technical difficulty with our internet provider and are working to resolve it as quickly as possible.

Friday, February 12, 2010


All phones at Dental are now working as they should be.

DENTAL CAMPUS - Phone outage

The Dental Campus is currently experiencing problems with it's phone system. The Patient Appointment Line(4-6700) is working but all else is non-functioning. We are currently working to resolve the issue and will update as we know more.

Tuesday, February 09, 2010

TitanConnect - 24/7

Despite the uncertainty related to the weather, the TitanConnect ( system will be open and available for use – 24/7.

Monday, February 08, 2010

TitanConnect E-mail Web Client Back Up

The TitanConnect web mail is back up and running.

TitanConnect E-mail Web Client Down

The web mail interface for TitanConnect is currently down. The Web Services Team is working to fix the issue.

Friday, February 05, 2010

All systems operational

At this time, all UDM systems appear to be operational.  If you experience any issues, please let us know via comment to this blog or Helpdesk ticket.